Lara needs 10 pounds of haddock for her dinner party. At the grocery store, Haddock potions weighing 4 3/8 + 2 1/10 are place on the scale.
How many pounds of Haddock are on the scale?
How many more pounds does Lara needs?
There are_____ pounds Haddock on the Scale?

Answer by stanbon(75887)

Lara needs 10 pounds of haddock for her dinner party. At the grocery store, Haddock potions weighing 4 3/8 + 2 1/10 are place on the scale.
How many pounds of Haddock are on the scale?
How many more pounds does Lara needs?
There are_____ pounds Haddock on the Scale?
--------------------------------------------------------
4 3/8 + 2 1/10
-----------------
= (8*4+3)/8 + (10*2+1)/10
-----------------
= 35/8 + 21/10
--------
= (10*35 + 8*21)/(80)
--------
= (350 + 168)/80
---
= 518/80
---
= 6.475 lbs of Haddock are on the Scale
----
She needs 10-6.475 = 3.5250 lbs. to meet her goal of 10 lbs.
====================
Cheers,
Stan H.
=================